A company is due to receive a payment Ksh 300,000 from a customer in 3 months time. To smooth it's cash flows, the company would prefer to receive the payment immediately an has agreed to transfer it's entitlement to this payment to a third party (a discount house)in return for an immediate payment calculated using a rate of commercial discount of 16%p.a. How much will the immediate payment made by the discount house be?
1
Expert's answer
2020-07-28T18:20:32-0400

I=Port

A=Po +I=PO+ Port=Po(1+rt)

·        I is the interest

·        A is the end amount principal plus interest

·        PO  principal

·        r is the interest rate in decimal form

·        t is time

A=Po +I=PO+ Port=Po(1+rt)

  =300000(1+16÷12/100×3)

    =300000(1+0.013×3)

     =300000×1.039

       =Ksh.311700